Grade 11IIT JEE Entrance Exam

100ml of 20.8% BaCl2 solution and 50ml of 9.8% H2So4 sol will form Baso4 -

SF
9 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
100ml of 20.8% BaCl2 solution ≡ 20.8 g BaCl2
50ml of 9.8% H2SO4 solution ≡ 4.9 g H2SO4
BaCl2 + H2SO4 = BaSO4 + 2HCl
208 g 98 g 233 g
98 g H2SO4 reacts with 208 g BaCl2
1 g H2SO4 reacts with 208/98 g BaCl2
4.9 g H2SO4 reacts with 208/98 X 4.9 g BaCl2 = 10.4 g BaCl2
98 g H2SO4 will produce 233 g BaSO4
1 g H2SO4 will produce 233/98 g BaSO4
4.9 g H2SO4 will produce 233/98 X 4.9 g BaSO4 = 11.65 g BaSO4
3.9